# Multi-modal AI for comprehensive breast cancer prognostication

### v1

@@ Abstract
Treatment selection in breast cancer is guided by molecular subtypes and clinical characteristics. Recurrence risk assessment plays a crucial role in personalizing treatment. Current methods, including genomic assays, have limited accuracy and clinical utility, leading to suboptimal decisions for many patients. We developed a test for breast cancer patient stratification based on digital pathology and clinical characteristics using novel AI methods. Specifically, we utilized a vision transformer-based pan-cancer foundation model trained with self-supervised learning to extract features from digitized H&E-stained slides. These features were integrated with clinical data to form a multi-modal AI test predicting cancer recurrence and death. The test was developed and evaluated using data from a total of 8,161 breast cancer patients across 15 cohorts originating from seven countries. Of these, 3,502 patients from five cohorts were used exclusively for evaluation, while the remaining patients were used for training. Our test accurately predicted our primary endpoint, disease-free interval, in the five external cohorts (C-index: 0.71 [0.68-0.75], HR: 3.63 [3.02-4.37, p<0.01]). In a direct comparison (N= 858), the AI test was more accurate than Oncotype DX, the standard-of-care 21-gene assay, with a C-index of 0.67 [0.61-0.74] versus 0.61 [0.49-0.73], respectively. Additionally, the AI test added independent information to Oncotype DX in a multivariate analysis (HR: 3.11 [1.91-5.09, p<0.01)]). The test demonstrated robust accuracy across all major breast cancer subtypes, including TNBC (C-index: 0.71 [0.62-0.81], HR: 3.81 [2.35-6.17, p=0.02]), where no diagnostic tools are currently recommended by clinical guidelines. These results suggest that our AI test can improve accuracy, extend applicability to a wider range of patients, and enhance access to treatment selection tools.
@@
Figure 1: We present a multi-modal AI test for invasive breast cancer. a, The key component of the test is a system which processes high-resolution digital images of breast cancer specimens. Features are extracted from the digitized slides using Kestrel, a foundation model trained using self-supervised learning on a pan-cancer dataset of 400 million pathology image patches. b, Extracted pathology features and clinical features are used to train supervised time-to-event models predicting breast cancer recurrence or death. c, The AI test produces a multi-modal risk score, integrating pathological and clinical risk scores. d, We developed the AI test using 4,659 patients across 10 cohorts from six countries, and evaluated it on 3,502 patients from five patient cohorts which were not used during training. Evaluation sets consisted of two cohorts (Providence and TCGA) with all invasive breast cancer subtypes, and three cohorts (UChicago, Basel, and Karmanos) with only HR+ HER2− patients tested with Oncotype DX. e, Recurrence rate increases monotonically as the AI test score increases. f, Patients with predicted high risk had significantly worse outcomes compared to patients with predicted low risk. g, The AI test performed consistently across all evaluation datasets. h, In the direct comparison to a standard-of-care genomic assay, Oncotype DX, the AI test was a better predictor of cancer recurrence. i, The AI test performs consistently across various endpoints.
@@ 1 Main
Over the past few decades, breast cancer treatment has evolved significantly with the introduction of various chemo, endocrine, and targeted therapies. Treatment selection is primarily guided by clinical guidelines that rely on clinicopathological factors such as tumor size, nodal status, and estrogen and HER2 receptor status. While this approach optimizes outcomes at the population level to a certain degree, it overlooks critical, actionable information for individual patients, exposing a significant gap in personalized care.
@@
Complementary to genomic assays, histopathological features are routinely evaluated by pathologists to stratify patient risk. These features include histologic grade, tumor infiltrating lymphocytes level, mitotic count, and Ki67 expression. It has been shown that these primarily morphological features add independent prognostic information to genomic scores, and recent prognostic models integrate both [6, 7]. However, these features are limited in their scope and prognostic ability, driving the need for stronger and more robust prognostic features. In recent years, advances in AI, particularly in self-supervised learning, have enabled the development of more effective methods for learning meaningful features from imaging data [8, 9, 10, 11, 12, 13]. These features are not based on pre-specified characteristics and extracting them does not require manual annotation by pathologists. Instead, AI models autonomously determine the most salient features by learning from millions of images. Early evidence shows that these AI-enabled features extracted from digital pathology images can be used to predict key patient characteristics and their long-term outcomes [14, 15, 16].
In response to the gaps in today’s treatment selection tools, we developed and evaluated an AI test for the stratification of invasive breast cancer patients. It is built upon the latest advances in self-supervised learning and digital pathology. Specifically, our test extracts strongly predictive morphological features from standard H&E-stained slides of core needle biopsy or surgical specimens utilizing Kestrel [17], a state-of-the-art pan-cancer AI foundation model. Kestrel is a variant of vision transformer [18] trained with self-supervised learning method DINOv2 [13] using 400M digital pathology image patches. The image patches used to train Kestrel were sourced from a large, heterogeneous database of pan-cancer whole slide images, including breast cancer images. Additionally, we tuned the hyperparameters for Kestrel using a novel search methodology that evaluates model performance on a suite of digital pathology benchmarks while iteratively scaling model and dataset size, yielding a model that is specially tuned for downstream tasks in digital pathology.
In this study, we present a multi-modal AI test that integrates pathology-based features, extracted by Kestrel, with routine clinical characteristics, namely cancer T and N stage, patient age, ER, PR and HER2 status and presence of ductal or lobular histology (Figure 1). The test produces a continuous score between 0 and 1 that is predictive of patient risk of cancer recurrence. The test was developed and evaluated using data from a total of 8,161 patients across 15 cohorts originating from seven countries. Of these, 3,502 patients from five cohorts were used exclusively for external evaluation, while the remaining patients were used for model training. Detailed patient characteristics are described in Table 1.
@@ 2 Results
Unless explicitly stated otherwise, the presented results are from the multi-modal AI test. We first report its prognostic capability, i.e., whether our model accurately predicts the risk of cancer recurrence and distinguishes between high- and low-risk patients. Then, we compare its accuracy against a standard-of-care genomic risk signature in various molecular, histological and clinical subgroups. Finally, we discuss potential clinical use cases. When reporting performance across multiple datasets, we present pooled results, using a random effects model (see Appendix A.7 and A.6 for performances on individual datasets).
To evaluate how well the test orders patients according to their risk, we report the concordance index (C-index), a metric analogous to AUROC which can be computed for censored outcomes. Additionally, we estimate the hazard ratio (HR) for the AI test in univariate and multivariate Cox proportional hazards models. The hazard ratio is estimated from the continuous score for every 0.2 unit increase in our score. Unless explicitly stated otherwise, we use the continuous AI test risk score, rather than its dichotomized version. Refer to Methods (Section 4) for a detailed explanation of the performance metrics used in this paper.
@@ 2.1 The AI test is predictive of breast cancer recurrence and survival
The AI test was found to be prognostic in all five patient cohorts used as external evaluation sets. For the primary endpoint, disease-free interval (DFI), the risk score generated by our model achieved a pooled C-index of 0.71 [0.68-0.75] and a pooled HR of 3.63 [3.02-4.37, p<0.01.
Two of the test datasets, Providence and TCGA-BRCA, represent a broad spectrum of invasive breast cancer patients. For prediction of DFI, our AI test achieved a C-index of 0.74 [0.70-0.79] and a HR 4.02 [3.09-5.23, p<0.01] in the Providence cohort (N= 1,733), and a C-index of 0.70 [0.63-0.77]] and a HR of 3.00 [2.10-4.28, p<0.01] in the TCGA-BRCA cohort (N=911).
Three cohorts — Karmanos, Basel, and UChicago — included only HR+ HER2− patients who were previously tested with Oncotype DX. The AI test achieved a C-index of 0.62 [0.49-0.75] and a HR of 3.82 [1.33-10.98, p=0.01] in the Karmanos cohort (N= 168), a C-index of 0.70 [0.60-0.80] and a HR of 3.98 [1.92-8.25, p<0.01] in the Basel cohort (N= 269), and a C-index of 0.67 [0.58-0.77] and a HR of 3.25 [1.45-7.31, p<0.01] in the UChicago cohort (N= 421).
The AI test was also prognostic for secondary endpoints. In analyses utilizing random effects models, pooled C-indices and HRs were as follows. For distant-disease free interval (DRFI), the test achieved a C-index of 0.70 [0.61-0.78] and a HR of 4.02 [2.64-6.13, p<0.01. For overall survival (OS), the test achieved a C-index of 0.65 [0.58-0.72] and a HR of 2.52 [1.88-3.38, p<0.01]. For recurrence-free survival (RFS), the test achieved a C-index of 0.68 [0.65-0.71] and a HR of 2.65 [2.33-3.01, p<0.01]. And for distant recurrence-free survival (DRFS), the test achieved a C-index of 0.66 [0.61-0.71] and a HR of 2.64 [2.14-3.26, p<0.01]. Forest plots for all endpoints and metrics are reported in Appendix A.6, and endpoints are defined in Appendix A.4.
@@ 2.2 The AI test is more accurate than a genomic assay in predicting cancer recurrence in HR+ patients
Figure 2: The AI test reclassifies patients tested with a standard-of-care genomic assay into different risk categories. Oncotype DX, a standard-of-care 21-gene assay, classifies patients into low-, intermediate-, and high-risk groups. The AI test demonstrated statistically significant discrimination between high- and low-risk patients without the need to introduce an intermediate-risk group, thereby enhancing clarity in decision-making. We analyze the differences in classification and patient outcomes between the two tests using data from the Karmanos, Basel, and UChicago cohorts. a, Comparison of prognostic ability between Oncotype DX and the AI test in univariate models. b, A scatter plot illustrating risk scores for Oncotype DX-tested patients. Each point represents one patient. The majority of patients with intermediate Oncotype DX scores were reclassified into the low-risk group by the AI test. c, For intermediate-risk Oncotype DX patients, the AI test was able to accurately distinguish between low- and high-risk patients. d, The AI test is significantly associated with a disease-free interval after adjusting for Oncotype DX score, grade and race in a multivariate Cox regression model. e, Comparing the AI test’s modalities in a multivariate Cox model shows that the pathology score was more informative than the clinical score within the Oncotype DX cohort.
@@
We compared the prognostic performance and potential clinical utility of our model to Oncotype DX, a standard-of-care 21-gene assay developed to predict distant recurrence risk. Three external cohorts (Karmanos, Basel, UChicago) contain a total of 858 patients who were tested with Oncotype DX Recurrence Score assays performed as part of routine care. Full results are in Appendix A.8.
@@
In the Karmanos cohort, the AI model’s C-index for DFI was 0.62 [0.49-0.76] and the HR was 3.81 [1.33-10.98, p<0.01], compared with Oncotype DX’s C-index of 0.54 [0.41-0.68] and HR of 1.36 [0.56-3.27, p=0.50]. In the Basel cohort, the AI test’s C-index was 0.70 [0.60-0.80] and the HR was 3.37 [1.54-7.40, p<0.01], compared with Oncotype DX’s C-index of 0.55 [0.42-0.68] and HR of 1.76 [0.85-3.64, p=0.13]. In the UChicago cohort, the AI test’s C-index was 0.67 [0.58-0.77] and the HR was 3.26 [1.45-7.31], compared with Oncotype DX’s C-index of 0.71 [0.61-0.81] and HR of 2.78 [1.61-4.81]. Together, the AI test achieved a pooled C-index of 0.67 [0.61-0.74] and HR of 3.67 [2.79-4.84, p<0.01], compared to Oncotype DX’s C-index of 0.608 [0.491-0.725] and HR of 2.09 [0.85-5.15, p=0.21]. These comparisons are illustrated in Figure 2A.
@@
Oncotype DX has been shown to predict the benefit of adjuvant chemotherapy in HR+ patients. Patients with high Oncotype DX scores had significant chemotherapy benefits, while patients with low Oncotype DX scores did not [19]. Large randomized clinical trials, TAILORx [20] and RxPONDER [21], have shown that some patients with intermediate Oncotype DX scores might also safely avoid adjuvant chemotherapy. For patients tested with Oncotype DX, the AI model would re-classify 666 out of 858 (77.6%) patients into different risk categories (Figure 2B). All intermediate Oncotype-risk patients ( 61.3% of all Oncotype DX patients) would be reclassified as either low- or high-risk by our AI test. 423 out of 526 (80.4%) intermediate Oncotype-risk patients would be classified as low-risk patients and 103 of 526 (19.6%) as high-risk patients using the AI test. Notably, 103 high Oncotype-risk patients (12.0% of all Oncotype DX patients) would be reclassified as low-risk by the AI test. Within this intermediate Oncotype-risk group, the continuous AI test score was significantly associated with recurrence (HR: 3.45 [1.85-6.42, p<0.01]) in the Karmanos, Basel and UChicago cohorts after adjusting for the dataset (see Table 5).
@@
Figure 3: The AI test performs well in all major clinically relevant groups. Forest plots display our AI test’s performance across clinical, molecular, and demographic groups. Next to each group name we report the number of disease-free interval events and total number of patients. Subgroup performances were pooled across all evaluation cohorts using a random effects model.
@@
Finally, we demonstrate that the AI test provides independent prognostic value alongside Oncotype DX as well as other common clinical indicators. To do this, we incorporated the AI test score into a mutlivariate Cox proportional hazards model along with Oncotype DX score, age, Nottingham cancer grade, dataset and race. Dataset was accounted for with an indicator variable for Karmanos and Chicago, with Basel as the reference dataset. Race was accounted for with an indicator variable for Other/Unknown, Asian, and Black, with White as the reference group. After adjusting for these factors, the AI test achieved a HR of 3.11 [1.91-5.09, p<0.01]. In comparison, Oncotype DX achieved a HR of 1.47 [0.93-2.30, p=0.10], which is not statistically significant. The HR between Karmanos and Basel was 0.91 [0.35-2.41, p=0.85], which is not significant. The hazard ratio between Chicago and Basel was 0.30 (0.09-1.00, p=0.05). None of the hazard ratios associated with race were significant. The full results are in Table 6 in Appendix A.5.
@@
2.3 The AI test performs well in various clinically meaningful subgroups. Potential clinical utility.

### v2

@@ Abstract
Treatment selection in breast cancer is guided by molecular subtypes and clinical characteristics. However, current tools including genomic assays lack the accuracy required for optimal clinical decision-making. We developed a novel artificial intelligence (AI)-based approach that integrates digital pathology images with clinical data, providing a more robust and effective method for predicting the risk of cancer recurrence in breast cancer patients. Specifically, we utilized a vision transformer pan-cancer foundation model trained with self-supervised learning to extract features from digitized H&E-stained slides. These features were integrated with clinical data to form a multi-modal AI test predicting cancer recurrence and death. The test was developed and evaluated using data from a total of 8,161 female breast cancer patients across 15 cohorts originating from seven countries. Of these, 3,502 patients from five cohorts were used exclusively for evaluation, while the remaining patients were used for training. Our test accurately predicted our primary endpoint, disease-free interval, in the five evaluation cohorts (C-index: 0.71 [0.68-0.75], HR: 3.63 [3.02-4.37, p<0.001]). In a direct comparison (n=858), the AI test was more accurate than Oncotype DX, the standard-of-care 21-gene assay, achieving a C-index of 0.67 [0.61-0.74] versus 0.61 [0.49-0.73], respectively. Additionally, the AI test added independent prognostic information to Oncotype DX in a multivariate analysis (HR: 3.11 [1.91-5.09, p<0.001)]). The test demonstrated robust accuracy across major molecular breast cancer subtypes, including TNBC (C-index: 0.71 [0.62-0.81], HR: 3.81 [2.35-6.17, p=0.02]), where no diagnostic tools are currently recommended by clinical guidelines. These results suggest that our AI test improves upon the accuracy of existing prognostic tests, while being applicable to a wider range of patients.
@@
Figure 1: We present a multi-modal AI test for invasive breast cancer. a, The key component of the test is a system which processes high-resolution digital images of breast cancer specimens. Features are extracted from the digitized slides using Kestrel, a foundation model trained using self-supervised learning on a pan-cancer dataset of 400 million pathology image patches. b, Extracted pathology features and clinical features are used to train supervised time-to-event models predicting breast cancer recurrence or death. c, The AI test produces a multi-modal risk score, integrating pathological and clinical risk scores. d, We developed the AI test using 4,659 patients across 10 cohorts from six countries, and evaluated it on 3,502 patients from five patient cohorts which were not used during training. Evaluation sets consisted of two cohorts (Providence and TCGA) with all invasive breast cancer subtypes, and three cohorts (UChicago, Basel, and Karmanos) with only HR+ HER2− patients tested with Oncotype DX. e, The 10 year recurrence probability increases monotonically as the AI test score increases. f, Patients with predicted high risk had significantly worse outcomes compared to patients with predicted low risk. g, The AI test achieved strong prognostic results across all evaluation datasets. h, In a direct comparison (n=858) to a standard-of-care genomic assay, Oncotype DX, the AI test was a better predictor of cancer recurrence. i, The AI test performs consistently across various endpoints (DFI - disease-free interval, DRFI - distant recurrence-free interval, RFS - recurrence-free survival, DRFS - distant recurrence-free survival, OS overall survival). Endpoint definitions are in Appendix A.4.
@@ 1 Main
Over the past few decades, breast cancer treatment has evolved significantly with the introduction of various chemo, endocrine, and targeted therapies. Treatment selection is primarily informed by clinical guidelines that rely on clinicopathological factors such as tumor size, nodal status, and estrogen and HER2 receptor status. While this approach optimizes outcomes at the population level to a certain degree, it overlooks critical, actionable information for individual patients, exposing a significant gap in personalized care.
@@
Complementary to genomic assays, histopathological features are routinely evaluated by pathologists to stratify patient risk. These features include histologic grade, tumor infiltrating lymphocytes level, mitotic count, and Ki67 expression. These primarily morphological features have been shown to add independent prognostic information to genomic scores, and recent prognostic models integrate both [6, 7]. However, these features are limited in their scope and prognostic ability, which drives the need for stronger and more informative prognostic features. In recent years, advances in AI, particularly in self-supervised learning, have enabled the development of more effective methods for learning meaningful features from imaging data [8, 9, 10, 11, 12, 13]. These features are not based on pre-specified characteristics and extracting them does not require manual annotation by pathologists. Instead, AI models autonomously determine the most salient features by learning from millions of images. Early evidence shows that these AI-enabled features extracted from digital pathology images can be used to predict key patient characteristics and long-term outcomes [14, 15, 16].
In response to the gaps in today’s treatment selection tools, we developed and evaluated an AI test for predicting the risk of cancer recurrence in invasive breast cancer patients. It is built upon the latest advances in self-supervised learning and digital pathology. Specifically, our test extracts strongly generalizable morphological features from standard H&E-stained slides of core needle biopsy or surgical specimens utilizing Kestrel [17], a state-of-the-art pan-cancer AI foundation model. Kestrel is a variant of vision transformer [18] trained with the self-supervised learning method DINOv2 [13] using 400M digital pathology image patches. The image patches used to train Kestrel were sourced from a large, heterogeneous database of pan-cancer whole slide images, including breast cancer images. Additionally, we tuned Kestrel’s hyperparameters using a novel methodology that evaluates model performance on a suite of digital pathology benchmarks. This approach iteratively scales model and dataset size, yielding a model that is specially tuned for downstream tasks in digital pathology.
In this study, we present a multi-modal AI test that integrates pathology-based features, extracted by Kestrel, with routine clinical characteristics, namely cancer T and N stage, patient age, ER, PR and HER2 status and presence of ductal or lobular histology (Figure 1). The test produces a continuous score between 0 and 1 that quantifies the risk of cancer recurrence. The test was developed and evaluated using data from a total of 8,161 patients across 15 cohorts originating from seven countries. Of these, 3,502 patients from five cohorts were used exclusively for external evaluation, while the remaining patients were used for model training. Detailed patient characteristics are described in Table 1.
@@ 2 Results
2.1 Breast cancer cohorts
To execute this project we assembled a diverse collection of both public datasets (TCGA-BRCA [19, 20], METABRIC [21], BASIS [22], Providence  [23]) and private datasets from several sources (NYU Langone Health, The Catholic University of Korea, Gundersen Health, National Center of Pathology in Lithuania, Breast Cancer Now Biobank, Cancer Research Malaysia, Omica.bio Cancer Atlas, Wales Cancer Biobank, Karmanos Cancer Institute, University Hospital Basel, and UChicago Medicine). Altogether, we collected 5162 slides from 4659 breast cancer patients across 10 datasets for training, and 3632 slides from 3502 breast cancer patients from 5 datasets for evaluation, making this one of the largest data collection efforts for predicting breast cancer recurrence. All patients in data are women. Two of the evaluation datasets, Providence (n=1733) and TCGA-BRCA (n=911), represent a broad spectrum of invasive breast cancer patients. The remaining three test cohorts — Karmanos (n=168), Basel (n=269), and UChicago (n=421) — included only HR+ HER2− patients who were previously tested with Oncotype DX. See Table 1 for more detailed statistics of the training and evaluation cohorts.
2.2 Study design and results
Unless explicitly stated otherwise, the presented results are from the multi-modal AI test. We first report its prognostic capability, i.e., whether our model accurately predicts the risk of cancer recurrence and distinguishes between high- and low-risk patients. Then, we compare its accuracy against a standard-of-care genomic risk signature in various molecular, histological and clinical subgroups. Finally, we discuss potential clinical use cases. When reporting performance across multiple datasets, we present pooled results using a random effects model. The additional results for specific data sets are in Figure A.10 and Figure A.11.
To evaluate how well the test orders patients according to their risk, we report the concordance index (C-index). Additionally, we estimate the hazard ratio (HR) for the AI test in univariate and multivariate Cox proportional hazards models. While the AI test takes values between 0 and 1, the HR presented is per every 0.2 unit increase. Unless explicitly stated otherwise, we use the continuous AI test risk score, rather than its dichotomized version. To compare hazard ratio with Oncotype DX, we dichotimize the score from the AI test. We defined high- and low-risk groups using the 80th percentile cutoff in the HR+ HER2- subpopulation within our largest internal training dataset (NYU). We refer to Methods (Section 4) for a detailed explanation of the performance metrics used in this paper.
@@ 2.3 The AI test is predictive of breast cancer recurrence and survival
The AI test was found to be prognostic in all five patient cohorts used as external evaluation sets. For the primary endpoint, disease-free interval (DFI), the risk score generated by our model achieved a pooled C-index of 0.71 [0.68-0.75] and a pooled HR of 3.63 [3.02-4.37, p<0.001].
Two cohorts — Providence and TCGA-BRCA — were composed of broad population of invasive breast cancer patients. Our AI test achieved a C-index of 0.74 [0.70-0.79] and a HR 4.02 [3.09-5.23, p<0.001] in the Providence cohort, and a C-index of 0.70 [0.63-0.77] and a HR of 3.00 [2.10-4.28, p<0.001] in the TCGA-BRCA cohort.
Three cohorts — Karmanos, Basel, and UChicago — included only HR+ HER2− patients who were previously tested with Oncotype DX. The AI test achieved a C-index of 0.62 [0.49-0.75] and a HR of 3.82 [1.33-10.98, p=0.013] in the Karmanos cohort (n=168), a C-index of 0.70 [0.60-0.80] and a HR of 3.98 [1.92-8.25, p<0.001] in the Basel cohort (n=269), and a C-index of 0.67 [0.58-0.77] and a HR of 3.26 [1.45-7.31, p=0.004] in the UChicago cohort (n=421).
The AI test was also prognostic for secondary endpoints (as defined in Appendix A.4). For distant-disease free interval (DRFI), the test achieved a C-index of 0.70 [0.61-0.78] and a HR of 4.02 [2.64-6.13, p=0.001]. For overall survival (OS), the test achieved a C-index of 0.65 [0.58-0.72] and a HR of 2.52 [1.88-3.38, p=0.001]. For recurrence-free survival (RFS), the test achieved a C-index of 0.68 [0.65-0.71] and a HR of 2.65 [2.33-3.01, p<0.001]. For distant recurrence-free survival (DRFS), the test achieved a C-index of 0.66 [0.61-0.71] and a HR of 2.64 [2.14-3.26, p<0.001]. Forest plots for all endpoints and metrics are reported in Figure A.8 and Figure A.9.
@@ 2.4 The AI test is more accurate than a genomic assay in predicting cancer recurrence in HR+ patients
Figure 2: The AI test enhances the prognostic accuracy over standard-of-care genomic assays for predicting cancer recurrence. Oncotype DX, a standard-of-care 21-gene assay, classifies patients into low-, intermediate-, and high-risk groups. The AI test demonstrated statistically significant discrimination between high- and low-risk patients without the need to introduce an intermediate-risk group, thereby enhancing clarity in decision-making. We analyze the differences in classification and patient outcomes between the two tests using data from the Karmanos, Basel, and UChicago cohorts. a, Comparison of prognostic ability between Oncotype DX and the AI test in univariate models. We compare the hazard ratio for every 0.2 increase in our score, compared to a 20 point increase for Oncotype DX. b, A scatter plot illustrating risk scores for Oncotype DX-tested patients. Each point represents one patient. The majority of patients with intermediate Oncotype DX scores were reclassified into the low-risk group by the AI test. c, For intermediate-risk Oncotype DX patients, the AI test was able to accurately distinguish between low- and high-risk patients (HR 2.84 [1.47-5.49, p=0.002]) d, Hazard ratios associated with common clinical covariates in a multivariate Cox analysis, with and without including the AI test. The AI test is significantly associated with DFI after adjusting for Oncotype DX score, grade (based on the Nottingham grading system, categorized from 1 to 3) and race in a multivariate Cox regression model. e, Comparing the AI test’s modalities in a multivariate Cox model shows that the pathology score was more informative than the clinical score within the Oncotype DX cohort (see Table A.4).
@@
We compared the prognostic performance and potential clinical utility of our model to Oncotype DX, a standard-of-care 21-gene assay developed to predict distant recurrence risk. Three external cohorts (Karmanos, Basel, UChicago) contain a total of 858 patients who were tested with Oncotype DX Recurrence Score assays performed as part of routine care.
@@
In the Karmanos cohort, the AI test’s C-index for DFI was 0.62 [0.49-0.76] and the HR was 3.82 [1.33-10.98, p=0.013], compared with Oncotype DX’s C-index of 0.54 [0.41-0.68] and HR of 1.36 [0.56-3.27, p=0.500]. In the Basel cohort, the AI test’s C-index was 0.70 [0.60-0.80] and the HR was 3.98 [1.92-8.25, p<0.001], compared with Oncotype DX’s C-index of 0.55 [0.42-0.68] and HR of 1.76 [0.85-3.64, p=0.13]. In the UChicago cohort, the AI test’s C-index was 0.67 [0.58-0.77] and the HR was 3.26 [1.45-7.31, p=0.004], compared with Oncotype DX’s C-index of 0.71 [0.61-0.81] and HR of 2.78 [1.61-4.81, p<0.001]. Together, the AI test achieved a pooled C-index of 0.67 [0.61-0.74] and HR of 3.67 [2.79-4.84, p=0.002], compared to Oncotype DX’s C-index of 0.608 [0.491-0.725] and HR of 2.09 [0.85-5.15, p=0.21]. These comparisons are illustrated in Figure 2A and further extended in Figure A.12.
@@
Oncotype DX has been shown to predict the benefit of adjuvant chemotherapy in HR+ patients. Patients with high Oncotype DX scores had significant chemotherapy benefits, while patients with low Oncotype DX scores did not [24]. Large randomized clinical trials, TAILORx [25] and RxPONDER [26], have shown that some patients with intermediate Oncotype DX scores might also safely avoid adjuvant chemotherapy. For patients tested with Oncotype DX, the AI model would re-classify 666 out of 858 (77.6%) patients into different risk categories (Figure 2B). All intermediate Oncotype-risk patients (61.3% of all Oncotype DX patients) would be reclassified as either low- or high-risk by our AI test. 423 out of 526 (80.4%) intermediate Oncotype-risk patients would be classified as low-risk patients and 103 of 526 (19.6%) as high-risk patients using the AI test. The outcomes for high/low risk patients are illustrated in Figure 2C. Within this intermediate Oncotype-risk group, the continuous AI test score was significantly associated with recurrence (HR: 3.45 [1.85-6.42, p<0.001]) in the Karmanos, Basel and UChicago cohorts after adjusting for the dataset (see Table A.5).
@@
Figure 3: The AI test performs well in all major clinically relevant groups. Forest plots display our AI test’s performance across clinical, molecular, and demographic groups. Next to each group name we report the number of disease-free interval events and total number of patients. Subgroup performances were pooled across all evaluation cohorts using a random effects model. For subgroup performance based on histological subtype, we excluded patients from the Providence cohort because histological subtype data were unavailable.
@@
Finally, we demonstrate that the AI test provides independent prognostic value from Oncotype DX, as well as from other common clinical indicators. To do this, we incorporated the AI test score into a multivariate Cox proportional hazards model along with Oncotype DX score, Nottingham cancer grade, dataset and race (see Section 4 for more details). After adjusting for these factors, the AI test has a HR of 3.11 [1.91-5.09, p<0.001]. In comparison, Oncotype DX has a HR of 1.47 [0.93-2.30, p=0.10], which is not statistically significant. Race groups or grade were not significant. Key findings are presented in Figure 2D. The full results are in Table A.6 in Appendix A.5. To further demonstrate the the AI test is not solely dependent on clinical covariates, we performed multivariate analysis with the clinical and pathology score within the AI test. As shown in Figure 2E, the pathology score is significant after adjusting for the clinical score. This shows that the pathology adds prognostic information that is independent from the information in clinical covariates used in our AI test.
@@
2.5 The AI test performs well in various clinically meaningful subgroups: potential clinical utility